Mott Run 2 Journal
- 1 Friday, October 23 DAY (Matt, Joe)
- 2 Friday, October 23 SWING (Tim, Marcy)
- 3 Saturday, October 24 OWL (Daniel, Riad)
- 4 Saturday, October 24 DAY (Marty, Joe)
- 5 Saturday, October 24 SWING (Tim, Matt)
- 6 Sunday, October 25 OWL (Daniel, Riad)
- 7 Sunday, October 25 DAY (Marty, Joe)
- 8 Sunday, October 26 SWING (Tim, Marcy)
Friday, October 23 DAY (Matt, Joe)
Started DAY shift with PSS down until ~1pm due to PSS BCM fault. Hall B beam at 31MHz recovered to FC1 and Vertical Wien filter set to +90 deg. Later INJ/NL locked-up and beam restored to FC2 fine. A "momentum measurement" setup was created, one with quads between IPM0L02-IPM0L03 nulled and minimal correctors; IPM GOF's are here. Central momentum of 5.487MeV/c was set using DL magnet. Crested gradients of 5.27MV/m(146.5deg) and 4.120MV/m(94.2deg) were found to be the best average of 2D line (no survey elements) and 5D line (incomplete survey information), expect both will be consistent at the <1%.
Friday, October 23 SWING (Tim, Marcy)
Started shift with white board tool-box meeting (Tim, Daniel, Matt, Marcy, Riad, Joe). The look up table of R028GSET values has an error - DO NOT use this table. We stepped through energy settings while keeping 0L launch fixed measuring a) 5D BPM's with fixed DL and then b) DL settings with fixed 5D BPM's. We later found the actual range of R028GSET was +/- 40keV, not +/- 400keV. Still the measurement is here.
Saturday, October 24 OWL (Daniel, Riad)
Beam was restored after bypass SL klystron. BCM0L02 cal factors had to be reset (ioc reboot?). Clean launch set to Mott, detector timing adjusted (DAQ or RF trigger drift?), PITA settings to null charge asymmetry corrected (orbit/transmission change since last calibration?), set vertical polarization and measured total rate vs. thickness. Detailed shift summary here.
Saturday, October 24 DAY (Marty, Joe)
During DAY shift we studied asymmetry vs. electron energy using the 1um and 0.35um gold foils about our nominal KE=5.0 MeV. We realized there was a typo in the energy look-up table, it spanned +/-40keV instead of +/-400keV so new gradients are being used and logged. We completed offsets of 0keV setup, -200keV setup, -400keV. At end of shift we are ~halfway done with +200keV setup and will then complete +400keV before completing this study. For most changes today the 0L02-8 auto-tuning was not working so well and converging so Shaun put it in manual mode and operation is fine there. We had two occasions of PSS BCM fault dropping the gun HV - so happening multiple times per shift.